A traditional individual retirement account (IRA) is a special type of retirement account in which the money you invest is exempt from income taxes until you withdraw it. If you deposit $100 each month into an IRA earning 6% interest, how much will you have in the account after 20 years?


1
Expert's answer
2021-10-06T17:20:43-0400

The monthly deposit "d= \$100"

6% annual rate "r=0.06"

12 months in 1 year "k=12"

We want the amount after 20 years "N=20"


Putting this into the equation:


"P_{20}=\frac{100
(
(
1
+\frac{
0.06}{
12}
)^{20
(
12
)}
−
1
)}
{(\frac{
0.06}{
12}
)}"


"P_{20}
=\frac{
100
(
(
1.005
)^{
240}
−
1
)}
{(
0.005
)}"


"P_{20}
=\frac{
100
(
3.310
−
1
)}
{(
0.005
)}"


"P_{20}
=\frac{
100
(
2.310
)}
{(
0.005
)}"

"P_{20}=\$
46200"


Professor P’s method:

We know the following: PMT=100, i=0.06/12=0.005, n=12(20)=240


"F
V
=\frac{
P
M
T
(
(
1
+
0.005
)^{
240}
−
1
)}
{(
0.005
)}=\$46,204.09"


The account will grow to $46,200 after 20 years.


Notice that you deposited into the account a total of $24,000 ($100 a month for 240 months). The difference between what you end up with and how much you put in is the interest earned.


In this case it is $46,204.09 – $24,000 = $22,204.09.